# Always Keep Your Promises: DynamicLRP, A Model-Agnostic Solution To Layer-Wise Relevance Propagation

### v1

@@ Abstract
Layer-wise Relevance Propagation (LRP) provides principled attribution for neural networks through conservation properties and foundations in Deep Taylor Decomposition. However, existing implementations operate at the module level, requiring architecture-specific propagation rules and model modifications. These limit the generality of target model, use cases in explanation-augmented training contexts, and sustainability of implementations as architectures evolve. We introduce DynamicLRP, a model-agnostic LRP framework operating at the tensor operation level. By decomposing attribution to individual operations within computation graphs and introducing a novel mechanism for deferred activation resolution, named the Promise System, our approach achieves true architecture agnosticity while maintaining LRP’s theoretical guarantees. This design operates independently of backpropagation machinery, enabling operation on arbitrary computation graphs without model modification and side-by-side execution with gradient backpropagation. Being based on computation graphs, this method is theoretically extensible to other deep learning libraries that support auto-differentiation. We demonstrate faithfulness matching or exceeding specialized implementations (1.77 vs 1.69 ABPC on VGG, equivalent performance on ViT, 93.70% and 95.06% top-1 attribution accuracy for explaining RoBERTa-large and Flan-T5-large answers on SQuADv2, respectively) while maintaining practical efficiency on models with hundreds of millions of parameters. We achieved 99.92% node coverage across 31,465 computation graph nodes from 15 diverse architectures, including state-space models (Mamba), audio transformers (Whisper), and multimodal systems (DePlot) without any model-specific code with rules for 47 fundamental operations implemented. Our operation-level decomposition and Promise System establish a sustainable, extensible foundation for LRP across evolving architectures.
Figure 1: ImageNette-320 example
Figure 2: Zennit LRP Attributions
Figure 3: Our LRP Attributions
@@ 1 Introduction
While significant progress has been made in explaining medium-scale models and established architectures, the decision-making processes of large-scale transformer systems Vaswani et al. (2017) remain challenging to interpret comprehensively. As these models increasingly influence high-stakes applications from healthcare to autonomous systems, understanding why they make specific predictions has become a critical scientific and societal imperative.
Rule-based decomposition methods, pioneered by Layer-wise Relevance Propagation (LRP) Bach et al. (2015), offer compelling results through systematic relevance redistribution. Unlike gradients, which reflect local sensitivity, LRP decomposes the actual prediction by backpropagating relevance scores while preserving conservation properties. This approach enables customizable propagation rules tailored to specific architectural components Montavon et al. (2017). This approach has proven remarkably effective for convolutional networks and demonstrated theoretical advantages over gradient-based alternatives Iwana et al. (2019); Böhle et al. (2019); Sun et al. (2021). However, LRP has historically required manual rule definition at each network layer, limiting its applicability to rapidly evolving architectures, where concepts like multi-head attention and residual connections demanded new specifications Achtibat et al. (2024); Otsuki et al. (2024).
This has led LRP to remain systematically underrepresented in contemporary explainability research, as it is one of the few mentioned methods which is not model-agnostic. Machiraju et al.’s Machiraju et al. (2024) comprehensive attribution study, which evaluated gradients, attention maps, SHAP, and LIME across transformer models, entirely omits LRP—despite its proven effectiveness and computational advantages.
We address this gap by proposing DynamicLRP, an LRP framework that extends rule-based attribution to arbitrary neural architectures with minimal overhead. We define LRP rules for individual tensor operations, analogous to generalized gradient computation found in auto-differentiation libraries (like PyTorch), which can be applied to any network topology. Through our novel "Promise System" for handling missing activations in the backward pass, our approach repurposes the computation graph used in auto-differentiation, and maintains LRP’s theoretical guarantees while achieving computational efficiency within the asymptotic bounds of standard backpropagation.
We aim to show through experiments across multiple modalities that our proposed LRP solution remains mathematically sound and produces faithful and interpretable attributions consistent with existing methods. We also demonstrate that the Promise-based LRP maintains a practical memory footprint and speed for research applications on medium-large (100s of millions - 1 billion parameters) models of various architectures using consumer-grade hardware.
We make three primary contributions:
  • Model-agnostic, operation-level LRP framework (Sec 4.1): We decompose LRP from module-level to operation-level, defining propagation rules for fundamental tensor operations rather than composite neural network modules to enable model agnosticity. Complete with rules for 47 fundamental tensor operations already implemented (See Appendix A).
  • Promise System (Sec 4.3): We introduce a novel mechanism for deferred relevance propagation and backward pass activation retrieval through the computation graph so that LRP can function without graph modification. The technique generalizes to arbitrary non-gradient computations on computation graphs used by auto-differentiation. We provide proven complexity bounds on memory and computation.
  • Faithfulness validation (Sec 5): We demonstrate faithfulness matching or exceeding specialized implementations across vision (ViT-b-16, VGG16) and text (RoBERTa, T5-Flan) while maintaining practical efficiency.
2 Background
2.1 Classic LRP via Taylor Approximation
Classic LRP redistributes the prediction score (relevance) backwards through the layers of a neural network. The key assumption is that the relevance at output neuron $j$ is proportional to the function value at $j$: $R_{j}\propto f_{j}(x)$. To propagate relevance to the inputs, we use the Taylor expansion and the conservation rule.
For a function $f_{j}(x)$, the first-order Taylor expansion around a root point $x_{0}$ is:
$f_{j}(x)\approx f_{j}(x_{0})+\sum_{i}\frac{\partial f_{j}}{\partial x_{i}}\Big|_{x_{0}}(x_{i}-x_{0,i})$ (1)
The conservation rule requires that the sum of input relevances equals the output relevance:
$\sum_{i}R_{i\leftarrow j}^{l-1}=R_{j}^{l}$ (2)
Assuming $R_{i\leftarrow j}^{l-1}$ is proportional to the Taylor term for input $i$:
$R_{i\leftarrow j}^{l-1}=\frac{\frac{\partial f_{j}}{\partial x_{i}}(x_{i}-x_{0,i})}{f_{j}(x)-f_{j}(x_{0})+\epsilon}R_{j}^{l}$ (3)
For a linear layer $z_{j}=\sum_{i}x_{i}w_{ji}+b_{j}$, and choosing $x_{0}=0$, this simplifies to:
$R_{i\leftarrow j}^{l-1}=\frac{x_{i}w_{ji}}{z_{j}+\epsilon}R_{j}^{l}$ (4)
where $\epsilon$ is a small stabilizer to avoid division by zero. The total relevance assigned to input $i$ is then $R_{i}=\sum_{j}R_{i\leftarrow j}$. This derivation shows how the epsilon rule arises from the proportionality assumption, Taylor expansion, and conservation principle, linking the mathematical definition of LRP to its practical implementation.
2.2 LRP Variants
2.2.1 $\gamma$-LRP
To denoise relevance attributions, the $\gamma$-LRP rule was introduced by Montavon et al. Montavon et al. (2019). It adds a hyperparameter $\gamma$ that acts as a multiplicative booster for positive neuron contributions:
2.2.2 Attn-LRP
Recent work by Achtibat et al. Achtibat et al. (2024) derived novel propagation rules using Deep Taylor Decomposition for both the softmax and bilinear matrix multiplication in the attention mechanism:
Bilinear Attention Matrix Multiplication
The relevance distribution for cell $A_{ji}$ in the matrix multiplication $O=A\cdot V$ is defined as:
$R_{ji}^{l-1}=\sum_{p}A_{ji}V_{ip}\frac{R_{jp}^{l}}{2O_{jp}+\epsilon}$ (6)
and $R_{ip}^{l-1}$ can be computed by summing over $j$ instead of $p$. This characterizes the distribution across the product of the attention logits and values matrices.
Softmax
The relevance at the input layer of the softmax is defined as:
$R^{l-1}=\sum_{i}x_{i}(R_{i}^{l}-s_{i}\sum_{j}R_{j}^{l})$ (7)
where $s_{i}$ is the softmax output and $x_{i}$ is the softmax input.
Our work builds on these insights to address the limitations of existing methods, providing a scalable and efficient solution for interpreting transformer models.
2.3 Computation Graphs in Deep Learning Frameworks
Deep learning frameworks rely on computation graphs to represent the sequence of operations performed on tensors and trace the gradients of the loss back to the network parameters. Formally, a computation graph is a directed acyclic graph (DAG) $G=(V,E)$, where $V$ is the set of nodes representing some operation $f_{v}$ (e.g., matrix multiplication, ReLU, addition), and $E$ is the set of directed edges representing data flow between operations.
In auto-differentiation libraries, each node $v\in V$ in the graph holds a function $\tilde{f}_{v}$ that computes the local gradient backpropagation of that operation using the Chain Rule, mapping the gradients of forward outputs to gradients of forward inputs. The edges $e\in E$ are thus directed from the model’s output layer to input layer, each vertex pointing to its forward dependencies. These functions must be well-defined for all tensor operations for the computation graph to be leveraged in backpropagation.
For example, a matrix multiplication node $v_{mm}$ from a linear layer will have $\tilde{f}_{mm}(\frac{\partial L}{\partial z})=\begin{pmatrix}\frac{\partial L}{\partial z}W^{T}&x^{T}\frac{\partial L}{\partial z}\end{pmatrix}$, where $L$ is the loss, $W$ is the weight matrix, $x$ is the layer input, and $z=Wx$. The in-edge pointing to $v$ would come from a node $v_{add}$, which represents the next operation in the forward pass, the bias addition $f_{add}(z,b)=z+b$. The two out-edges leaving $v$ point to a terminal node representing $W$, indicating it is a parameter, and the node representing the operation which produced $x$ in the forward pass.
Our proposed attribution algorithm extends this concept by associating each node in the graph with a custom propagation rule for relevance. This modular approach allows for operation-specific attributions, enabling the algorithm to handle complex architectures like transformers as standard functionality. By leveraging the DAG structure constructed by a reliable learning framework, the algorithm ensures that relevance is propagated efficiently and faithfully. The LRP conservation principle is thus adhered to on an operation-wise basis, instead of layer-wise.
$\sum_{i\in\text{inputs}}R_{i}=\sum_{j\in\text{outputs}}R_{j},$ (8)
where $R_{i}$ and $R_{j}$ are the relevance scores at the input and output nodes, respectively.
@@ 3.1 Explainability Methods
Explainable AI (XAI) algorithms fundamentally seek to decompose model predictions into input feature contributions, yet differ dramatically in their mathematical foundations and computational approaches. Gradient-based methods leverage the chain rule to compute $\frac{\partial f}{\partial x}$ as attribution signals Simonyan et al. (2014); Sundararajan et al. (2017); Smilkov et al. (2017), but suffer from gradient saturation, gradient shattering, and linearity assumptions that fail to capture complex nonlinear interactions Ancona et al. (2018); Samek et al. (2021). Perturbation-based approaches like SHAP Lundberg & Lee (2017) and LIME Ribeiro et al. (2016) directly measure prediction sensitivity through game-theoretic frameworks or local surrogates, yet incur exponential computational costs, making them intractable for high-dimensional data Covert et al. (2020). Class Activation Mapping methods Zhou et al. (2016); Selvaraju et al. (2017) leverage architectural constraints for spatial localization but provide coarse-grained explanations lacking principled mathematical foundations for complex architectures. This positions LRP as a strong explainability method when used in the correct domains, as it does not suffer from these pitfalls due to its tight rooting in the model architecture and efficient computation requirements of a single backward pass Samek et al. (2021).
3.2 LRP Implementations
We examine and compare against Zennit, an explainability library built for PyTorch and focused on LRP and other rule-based methods Anders et al. (2021). As of currently, Zennit is the most flexible, readily available, and up-to-date implementation of LRP, when compared to other libraries such as Captum Kokhlikyan et al. (2020) and iNNvestigate Alber et al. (2018), which both make LRP customization difficult Anders et al. (2021). Anders et al. base their framework on the PyTorch Module structure, making great use of Module Hooks to transform the Autograd backward pass to propagate relevance instead of gradients. They also introduce multiple user-facing APIs (Module rules, Composites, and Canonizers) to enable the mapping of foundational LRP rules to new modules and architectural patterns. However, adapting these to unseen patterns is a task in and of itself.
AttnLRP, a recent advancement which derived new propagation rules suitable for attention modules, provides an implementation named LxT (LRP eXplains Transformers) built on top of Zennit which highlights the limitations of this model-to-model configuration-based design Achtibat et al. (2024). The list of architectures currently supported by LxT is limited to Llama 2/3, Gemma 3, Qwen 2/3, BERT, GPT-2, and Vision Transformers. Each model requires a custom copied implementation of itself which patches and hooks the Attention, Linear, Convolution, Normalization, and Activation modules to new functions which invoke the LRP propagation rules for those modules.
While implementation of new rules is claimed to be trivial in Zennit, there is undoubtedly reasonable conceptual overhead to understanding how to correctly merge, split, and modify modules to fit within the framework, and additional management overhead with the requirement for new configurations and codebase semantics for adding support for other architectures.
@@ 4.1 Operation-Level Relevance Propagation
We motivate shifting the level of abstraction of LRP from layer/module-wise to operation-wise through careful consideration of the modern-day deep learning paradigm and the practices used by existing LRP implementations. We define a module as an object lying at the maximal granularity within a model structure defined by a deep learning framework (e.g. Linear, Convolution, Attention, LayerNorm). While many modules are defined by a sole operation such as Linear, Convolution, Softmax, etc., newer modules compound many elementary tensor operations, which are not captured at the granularity of modules.
The first consequence is that there is no upper limit for the number of unique module types, each requiring a new backpropagation rule. This is unsustainable in the long-term. By shifting to the level of operations, there is a strict set of operation-specific rules that must be defined before all architectures that reduce to that set of operations are supported.
Figure 4: Visualizing the difference between a module-level and operation-level approach for a simplified Transformer block. The rightmost diagram represents some fabricated toy example architecture that shares the same operation set as the Transformer, and thus does not require any new implementation to be covered by operation-level LRP, but would need new rules and configurations under module-level LRP.
Another complication is that operations not captured at the module level must always appear only as sub-steps within a module-specific propagation rule if they are to be considered in LRP. Thus, redundancy arises when multiple modules make use of the same elementary operations, but still require two unique propagation rule definitions.
Our operation-level approach modularizes relevance propagation, reducing any model architecture to a fixed and bounded set of fundamental tensor operations, treating operations as first-class citizens in propagation. We implement propagation rules for each type of Node in the computation graph (a Node in this case maps to an operation call).
4.2 Propagating Through the Computation Graph
We traverse the computation graph with the following heuristic:
Definition 1 (Traversal Heuristic).
Given some DAG $G=(V,E)$, for all $v\in V$, let $pending(v)=indegree(v)$. Let outadjoutadj be the graph defined as an out-adjacency list.
Initialize $S$ as a stack. Push all nodes with in-degree 0 onto $S$. Traverse as follows:
  • Pop the top node $v$ from the stack and process it (in our case, apply the operation-specific propagation function).
  • For all $w\in outadj(v)$, decrement $pending(w)$ by 1.
  • For all $w^{\prime}\in outadj(v)$ where $pending(w^{\prime})==0$ now, push $w^{\prime}$ onto $S$.
  • Repeat while $S$ is non-empty.
This ensures traversal follows some topological ordering of $G$, since nodes are only processed once all their in-neighbours have been.
While the above solution seems sound, a subtle issue lies within defining the propagation functions for each Node type. Consider a simple approach for backpropagating $R_{c}$ to $R_{a}$ and $R_{b}$ for $c=a+b$ as follows using absolute value ratios:
$R_{a}=R_{c}\times\frac{a.abs()}{a.abs()+b.abs()}\quad\text{and}\quad R_{b}=R_{c}\times\frac{b.abs()}{a.abs()+b.abs()}$ (9)
We suggest this because we want to attribute relevance to each tensor component proportionally to the magnitude of its contribution to the operation result.
Since computation graphs are mainly used for auto-differentiation, each Node only caches information necessary for computing the gradients w.r.t. its arguments.
But for $c=a+b$, we see that $\frac{\partial c}{\partial a}=\frac{\partial c}{\partial b}=1$. This implies that $a$ and $b$ need not be stored in any Addition Node, which indeed is the case in a framework such as PyTorch. This poses a critical problem for us, since without $a$ and $b$, we are then unable to compute $R_{a}$ and $R_{b}$ faithfully when we traverse an Addition Node, and this operation is not the only example of this case.

### v3

# Always Keep Your Promises: A Model-Agnostic Attribution Algorithm for Neural Networks

@@ Abstract
Inability to precisely understand neural network outputs is one of the most severe issues limiting the use of AI in multiple domains, from science and medicine to high-stakes decision and regulatory models. Layer-wise Relevance Propagation (LRP) is an established explainability method that addresses some of these limitations, but widespread adoption has not been possible because existing implementations must be coupled with individual model architectures, rendering them impractical for a rapidly evolving model space. We present DynamicLRP, a lightweight and flexible method for performing LRP on any neural network with provably minimal overhead. To achieve this, we introduce a novel graph search mechanism called the “Promise System”, which repurposes deep learning computation graphs for non-gradient computations and is implemented at the level of primitive tensor operations using standard automatic differentiation libraries. We demonstrate that DynamicLRP matches or surpasses specialised implementations in attribution quality across vision and language tasks and remains efficient for models at the billion-parameter scale. Notably, our implementation achieved 99.99% operation coverage across 31,465 computation nodes from 15 diverse architectures, without any architecture-specific modifications. This is the first truly model-agnostic LRP solution, enabling high-quality neural network attribution across the full spectrum of modern AI architectures. All code is available at https://github.com/keeinlev/dynamicLRP.
@@ 1 Introduction
Modern neural networks, in particular massive transformer-based foundation models Gemma Team (2025); Dosovitskiy et al. (2021); Grattafiori et al. (2024), offer unprecedented ability to map complex, unstructured inputs to accurate outputs. Inability to understand this mapping severely limits use of these powerful models in all domains where the reverse mapping is wanted. Pertinent domains include, but are not limited to: natural science, social science, medicine, law, and finance.
While a variety of methods for Explainable AI (XAI) do exist, each has substantive drawbacks. The most well-known methods include perturbation-based approaches such as SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016). These methods are thoroughly unsuitable for use in massive models, with SHAP explanations becoming more misleading with input cardinality Huang and Marques-Silva (2024), and LIME suffering from an arbitrary definition of sensitivity to locality. These methods are further troubled by the exponential computation incurred proportional to model complexity Covert et al. (2020). Gradient-based methods Simonyan et al. (2014); Sundararajan et al. (2017); Smilkov et al. (2017) are fragile, suffering from saturation, shattering, and assumptions of linearity Ancona et al. (2018); Samek et al. (2021). Class Activation Mapping Zhou et al. (2016); Selvaraju et al. (2017) lacks granular precision. In contrast, Layer-wise Relevance Propagation (LRP), a backpropagation-based method, Bach et al. (2015) provides a principled, efficient alternative by decomposing predictions into input contributions via strict conservation properties Montavon et al. (2017), avoiding the pitfalls of local sensitivity Samek et al. (2021) and proving effective in diverse domains Iwana et al. (2019); Böhle et al. (2019); Sun et al. (2021).
Despite its theoretical advantages, LRP remains constrained by practical limitations tied to module-specific rules, severely limiting adaptability to rapidly evolving architectures like Transformers Achtibat et al. (2024); Otsuki et al. (2024). Modern libraries like iNNvestigate Alber et al. (2018), Captum Kokhlikyan et al. (2020), and notably Zennit Anders et al. (2021) attempt to mitigate this with automated differentiation, but remain module-centric: new architectures typically need new composite rules and custom configurations. Even specialized advances like AttnLRP Achtibat et al. (2024) (built on Zennit) require manual adaptation for new model updates (e.g., Llama 2 vs. 3), creating friction and overhead for continued use.
Figure 1: Relevance flow diagrams for sample transformer blocks demonstrate the contrast in design. Low-level implementation allows Dynamic LRP adaptation to any arrangement of covered operations, such as the toy picture on the right. Note that as relevance flows from output to input, the arrow direction opposes inference.
To address these limitations, we introduce DynamicLRP, which shifts the focus from a module-centric to a primitive-based formulation by redefining the propagation rules at the level of atomic tensor operations (e.g., addition, matrix multiplication). Analogous to generalized gradient computation, DynamicLRP implements LRP rules for 47 primitives and propagates relevance directly through the computation graph, making the method inherently model-agnostic and free of model-specific configuration (Figure 1). This design is enabled by the Promise System, a novel deferred propagation mechanism that resolves the challenge of missing operand information during non-gradient computations within auto-differentiation frameworks (Sec 4.2). Lastly, we show that our approach matches or exceeds specialized implementations in attribution quality (Sec 5) across popular portable vision (ViT-b-16 Dosovitskiy et al. (2021), VGG16 Simonyan and Zisserman (2015)) and language (LLaMA Grattafiori et al. (2024), BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), T5-Flan Lee (2023)) models while maintaining efficiency on models into the billion-parameter scale.
@@ 2.1 Explainability Methods
Explainable AI (XAI) algorithms fundamentally seek to decompose model predictions into input feature contributions, yet differ dramatically in their mathematical foundations and computational approaches. Gradient-based methods leverage the chain rule to compute $\frac{\partial f}{\partial x}$ as attribution signals Simonyan et al. (2014); Sundararajan et al. (2017); Smilkov et al. (2017), but suffer from gradient saturation, gradient shattering, and linearity assumptions that fail to capture complex nonlinear interactions Ancona et al. (2018); Samek et al. (2021). Perturbation-based approaches like SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016) directly measure prediction sensitivity through game-theoretic frameworks or local surrogates, yet incur exponential computational costs, making them intractable for high-dimensional data Covert et al. (2020). Class Activation Mapping methods Zhou et al. (2016); Selvaraju et al. (2017) leverage architectural constraints for spatial localization but provide coarse-grained explanations lacking principled mathematical foundations for complex architectures. This positions LRP as a strong explainability method when used in the correct domains, as it does not suffer from these pitfalls due to its tight rooting in the model architecture and efficient computation requirements of a single backward pass Samek et al. (2021).
2.2 Modern LRP Implementations
LRP has evolved significantly since its inception, with modern implementations striving to achieve greater flexibility and efficiency. Several libraries, including iNNvestigate Alber et al. (2018), Captum Kokhlikyan et al. (2020), and Zennit Anders et al. (2021), provide implementations of LRP alongside other interpretability methods. Among these, Zennit stands out for its comprehensive support for rule-based methods in PyTorch, offering a modular framework that leverages the automatic differentiation machinery to backpropagate relevances intead of gradients. Another recent advancement in the field that is close to our work is AttnLRP Achtibat et al. (2024), which derived new propagation rules suitable for attention modules and provides an implementation named LxT (LRP eXplains Transformers) built on top of Zennit. Note that at the time of writing, the list of architectures currently supported by LxT is limited to Llama 2/3, Gemma 3, Qwen 2/3, BERT, GPT-2, and Vision Transformers. While LxT demonstrates the potential of extending LRP to modern architectures, it also illustrates the challenges of adapting module-centric designs and highlights the importance of a more flexible, architecture-agnostic approach.
While the implementation of new rules is claimed to be trivial in Zennit, there is nonetheless significant conceptual overhead in understanding how to correctly merge, split, and modify modules to fit within the framework, as well as additional management overhead due to the need for new configurations and codebase semantics when supporting other architectures. In particular, each model requires a custom, duplicated implementation to integrate the appropriate LRP propagation rules throughout the network. We build upon the foundation established by LxT and Zennit, acknowledging their flexibility while addressing the limitations of its module-centric approach. Specifically, adapting Zennit to previously unseen architectural patterns often requires the manual definition of new module mappings, a task that our operation-level approach seeks to automate.
3 Preliminaries
3.1 Classic LRP via Taylor Approximation and LRP variants
The classic LRP formulation as presented by Bach et al. (2015) derives LRP using what is called the deep Taylor decomposition framework. This formulation is functional, treating the layers as mathematical operators, and thus it is devoid of the structural assumptions of any given architecture. By taking the first-order Taylor expansion of the function at some root point $x_{0}$ such that $f(x_{0})=0$, and invoking the proposed LRP relevance conservation property, they define a recursive expression for the relevance of neurons $x_{j}$ at layer $l$ w.r.t. neurons $x_{i}$ at layer l1l-1. For a detailed derivation, see Appendix A
The $\epsilon$-LRP Rule Assuming the bias contribution is negligible or handled separately (see Achtibat et al. (2024) for a detailed discussion on bias absorption), the input relevance can be computed by summing over all output neurons and introducing a stabilizing factor $\epsilon$ to properly handle vanishing denominators. For a linear layer $z_{j}=\sum_{i}x_{i}w_{ji}+b_{j}$ with $x_{0}=0$:
$R_{i\leftarrow j}^{l-1}=\frac{x_{i}w_{ji}}{z_{j}+\epsilon}R_{j}^{l}$ (1)
This recovers the standard $\epsilon$-LRP rule. In matrix form, efficiently implementable on GPUs:
$R^{l-1}=x\odot(W^{\top}\cdot(R^{l}\oslash(z+\epsilon)))$ (2)
[…]
The $\gamma$-LRP Rule Attributions can become noisy when using the $\epsilon$-LRP rule, particularly in scenarios where the denominator $z_{j}+\epsilon$ becomes dominated by the stabilizing factor $\epsilon$, resulting in locally insensitive relevance scores. The $\gamma$-LRP rule Montavon et al. (2019), can be used to address this by introducing the hyperparameter $\gamma$ to act as a multiplicative booster for positive neuron contributions:
The Deep Taylor Decomposition framework has been recently used to derive propagation rules for all components in transformer-based architectures Vaswani et al. (2017). In their formulation, Achtibat et al. Achtibat et al. (2024) successfully derive rules for both bilinear matrix multiplications and softmax operations used in attention mechanisms. Our work builds on these insights to enable the reuse of these primitives for novel architectures.
3.2 Computation Graphs in Deep Learning Frameworks
Our attribution algorithm relies on the Backward Computation Graph, used by most deep learning frameworks (e.g., PyTorch, TensorFlow) to facilitate automatic differentiation. Informally, the backward computation graph of a neural network is a Directed Acyclic Graph (DAG) $G=(V,E)$ whose nodes $V$ correspond to operations executed during the forward pass, and whose directed edges $(u,v)\in E$ indicate that the output of operation $v$ was used as an input to operation $u$ in the forward pass, so that edges are oriented from outputs toward inputs.
The in-degree and out-degree of a node denote the number of operations that consumed its output and the number of inputs it consumed, respectively. Hence, the model output serves as the source node i.e., with zero incoming edges or in-degree of 0, while model parameters and data inputs are sink nodes i.e., with zero outgoing edges or out-degree of 0. For any node $u$, its outgoing adjacency set $outadj(u)=\{v|(u,v)\in E\}$ represents the set of operations that produced the inputs for $u$ (i.e., its downstream dependencies in the backward pass), the converse being its incoming adjacency set $inadj(u)=\{v|(v,u)\in E\}$. Our method leverages this structure to propagate relevance scores in a manner analogous to gradient computation, ensuring that at each step, the necessary information is available for accurate attribution.
@@ 4 Methods
We motivate shifting the level of abstraction of LRP from layer/module-wise to operation-wise through careful consideration of the modern-day deep learning paradigm and the practices used by existing LRP implementations. We define a module as an object lying at the maximal granularity within the model structure of a deep learning framework (e.g. Linear, Convolution, Attention, LayerNorm). While many modules are defined by a single operation, such as Linear, Convolution, Softmax, etc., certain modules compound many elementary tensor operations, which are not captured at the granularity of modules. There is thus redundancy in redefining propagation steps for tensor primitives within different module composite rules. The other consequence is that there is no upper limit for the number of unique module types, and each requires a new backpropagation rule. This is unsustainable in the long-term. By shifting to the level of operations, the rules for a set of operations needs to only be defined once before all architectures (and thus modules) that decompose to that set of operations are supported.
4.1 Topological Propagation and the Problem of Missing Activations
Given the backward computation graph $G=(V,E)$ defined in Section 3.2, our goal is to propagate relevance from the source (model output) to the sinks (inputs/parameters). Note that this operation-wise view contrasts with the classic layer-wise view (Section 3.1) but is mathematically equivalent, as layers are simply subgraphs of operations. Conservation of relevance applies locally at each node (i.e. $\sum_{v\in outadj(u)}R_{u\to v}=\sum_{p\in parents(u)}R_{p\to u}$)
Ideally, LRP should traverse $G$ in a way such that a node $v$ processes relevance only after it has received all relevance contributions from its parents in $G$, so we must visit nodes in an order that respects dependencies. Specifically, a node $v$ should only compute and distribute its relevance once it has aggregated relevance from all nodes $u$ that depend on it (i.e., all $u\in inadj(v)$). This dependency constraint is satisfied by traversing $G$ in some topological order. We present the graph traversal algorithm used for this purpose in Algorithm 1.
Algorithm 1 Topological Traversal of Backward Graph
1:Input: Backward Graph $G=(V,E)$
2:Initialize $S$ as a stack containing all nodes with $indegree(v)=0$
3:Initialize $pending(v)=indegree(v)\quad\forall v\in V$
4:while $S$ is not empty do
5:  Pop node $v$ from $S$
6:  Process($v$) $\triangleright$ Apply operation-specific propagation
7:  for $w\in outadj(v)$ do
8:    $pending(w)\leftarrow pending(w)-1$
9:    if $pending(w)==0$ then
10:     Push $w$ onto $S$
11:    end if
12:  end for
13:end while
While the topological traversal ensures the correct order of operations, propagating the relevance will fail in some cases due to the information available at each node. Consider the propagation rule for addition $c=a+b$, which distributes the relevance proportional to the contribution magnitude:
$R_{a}=R_{c}\times|a|/(|a|+|b|)\quad\text{and}\quad R_{b}=R_{c}\times|b|/(|a|+|b|)$ (4)
To compute $R_{a}$ and $R_{b}$, we strictly require the values of operands $a$ and $b$. However, modern automatic differentiation frameworks optimize memory by discarding $a$ and $b$ after the forward pass since $\frac{\partial c}{\partial a}=\frac{\partial c}{\partial b}=1$. Thus, even if the topological traversal visits the Addition node at the correct time, the operand tensors required for the attribution rule are missing from the graph.